OpenStudy (anonymous):

The formula for degrees Celsius is: C =5/9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 20 degrees Celsius is.

OpenStudy (anonymous):

@AravindG help

OpenStudy (aravindg):

first multiply by 9/5 on both sides

OpenStudy (anonymous):

srry my pc was messing up ok

OpenStudy (aravindg):

ok did you do that ?

OpenStudy (anonymous):

so would it be 9/5*9/5 & (F-32)*9/5

OpenStudy (aravindg):

5/9 (F − 32) X 9/5

OpenStudy (anonymous):

this is where i get confused b/c of the fractions so how would i do the next step

OpenStudy (aravindg):

you can cancel off 9/5 and 5/9

OpenStudy (anonymous):

so that leaves me (F-32)X9/5

OpenStudy (aravindg):

nope that leaves u with (F-32)

OpenStudy (anonymous):

but you multiplied 9/5 X(F-32)

OpenStudy (aravindg):

5/9 (F − 32) X 9/5= (F-32)=C (9/5)

OpenStudy (anonymous):

so then what

OpenStudy (aravindg):

then add 32 on both sides

OpenStudy (ajprincess):

have u solved for F? @R00nnYBraiNsbiG

OpenStudy (anonymous):

idk that was a test that i took on friday

OpenStudy (ajprincess):

(F-32)=C (9/5) What do u get when u add 32 to both sides as @AravindG said?